OpenStudy (anonymous):

Evaluate the integral: 3/square root(4-9x^2)dx

sam (.sam.):

Same, use \[x=asin \theta\]

OpenStudy (mimi_x3):

\[\int\limits\frac{3}{\sqrt{4-9x^{2}}} dx => 3\int\limits\frac{1}{\sqrt{4-9x^{2}}} =>3\int\limits\frac{1}{\sqrt{9(\frac{4}{9}-x^{2})}} \] It's a trig inverse function..